Calculate the resultant vector for the following forces:


1)     N = 8.9/ Direction = 39.0˚

FR =

2)     N = 15.5/ Direction = -149.1˚

FR =

1
Expert's answer
2021-10-05T10:07:44-0400

"\alpha=360-39-149.1=171.9°,"

"F=\sqrt{F_2^2+F_2^2-2F_1F_2\cos\alpha}=24.3."